# Rewriting the Code: A Simple Method for Large Language Model Augmented Code Search

### v1

@@ Abstract
In code search, the Generation-Augmented Retrieval (GAR) framework, which generates exemplar code snippets to augment queries, has emerged as a promising strategy to address the principal challenge of modality misalignment between code snippets and natural language queries, particularly with the demonstrated code generation capabilities of Large Language Models (LLMs). Nevertheless, our preliminary investigations indicate that the improvements conferred by such an LLM-augmented framework are somewhat constrained. This limitation could potentially be ascribed to the fact that the generated codes, albeit functionally accurate, frequently display a pronounced stylistic deviation from the ground truth code in the codebase. In this paper, we extend the foundational GAR framework and propose a simple yet effective method that additionally Rewrites the Code (ReCo) within the codebase for style normalization. Experimental results demonstrate that ReCo significantly boosts retrieval accuracy across sparse (up to 35.7%), zero-shot dense (up to 27.6%), and fine-tuned dense (up to 23.6%) retrieval settings in diverse search scenarios. To further elucidate the advantages of ReCo and stimulate research in code style normalization, we introduce Code Style Similarity, the first metric tailored to quantify stylistic similarities in code. Notably, our empirical findings reveal the inadequacy of existing metrics in capturing stylistic nuances.
@@ 2 Related Works
Recently, transformer-based pre-trained models significantly outperformed previous methods, since they can be trained on large-scale unlabelled corpus with self-supervised pre-training tasks. Many novel pre-training tasks are proposed to let models have a better general understanding of codes Guo et al. (2021); Li et al. (2022b, c); Shi et al. (2022). For instance, CodeBERT Feng et al. (2020) utilizes masked language modeling and replaced token detection. CodeT5 Wang et al. (2021) focuses on generative tasks through bimodal dual generation. UniXcoder Guo et al. (2022) integrates the aforementioned generative and understanding pre-training tasks.
@@
While LLMs [^1] We introduce related works for LLMs in Appendix A. are designed for token generation, their direct application to retrieval tasks such as code search is not suitable. Indeed, there have been attempts to amalgamate the search query and all the candidates together as input, subsequently requesting the LLMs to rank the candidates within the input Qin et al. (2023). However, the constraint on input sequence length impedes its applicability to large-scale retrieval tasks.
One indirect way is to ask LLMs to generate some references and expand the search query with them. This framework, denoted as Generation-Augmented Retrieval, has been proven effective in both question answering and passage retrieval Mao et al. (2020); Gao et al. (2022); Wang et al. (2023). Mao et al. (2020) is the first work to propose GAR in question answering. HyDE Gao et al. (2022) evaluates GAR in passage retrieval under zero-shot setting. query2doc Wang et al. (2023) extends GAR to fine-tuning. Our research findings suggest that the Generation-Augmented Retrieval (GAR) method does not substantially enhance the efficiency of code search, primarily due to the significant stylistic difference between exemplar code and true code.
@@ 3.2 ReCo
Detailed examples are provided in Appendix B to further elucidate the prompt.
@@
We offer theoretical insights to more effectively differentiate between GAR and our proposed ReCo. Each code in the codebase is an implementation of a specific query, which we denote as $c\sim P(q)$. Here $P$ denotes a real-world distribution between queries and codes. LLM also defines a probability distribution over queries. Thus, exemplar codes can be considered to follow $c_{q}\sim\mathrm{LLM}(q)$. We could find that $c$ and $c_{q}$ are sampled from two different distributions given $q$. This accounts for the occasional divergence between the true code and the exemplar code for the same query, as illustrated in Fig. 1.
@@ 5 Experimental Setups
We evaluate ReCo across various search scenarios and programming languages: online forum StackOverflow CoNaLa Yin et al. (2018), coding challenge competence APPS Hendrycks et al. (2021), general programming problems MBPP Austin et al. (2021) and MBJP Athiwaratkun et al. (2022). The first three datasets are written in Python while the last one is written in Java. The statistics of the datasets are shown in Appendix C.1. We take the widely adopted Mean Reciprocal Rank (MRR) as the evaluation metric Li et al. (2023). MRR is the average of reciprocal ranks of the true code snippets for the given query.
@@
We apply ReCo on several models: BM25, an enhanced version of TF-IDF, is a statistical measure that matches certain keywords in codes with the given query. CodeBERT is a bi-modal model pre-trained on Masked Language Modeling and Replaced Token Detection. UniXcoder unifies both generating and understanding pre-training tasks to further enhance code representation learning by leveraging cross-modal contents. Contriever Izacard et al. (2021) is an unsupervised dense information retrieval model that leverages contrastive learning for its training.
Compared Metrics
Implementation Details
When prompting LLMs, we randomly sample 4 in-context examples from the training sets and set a temperature of 1. And when we prompt LLMs multiple times for the same input, each time we resample the in-context example. The maximum length of output for code summarization and generation is 128 and 256, respectively. For sparse retrieval, we use the default implementation from Pyserini Lin et al. (2021). For dense retrieval, during training, we adopt the default hyperparameters described in the original paper. They are trained for 10 epochs with a batch size of 32. Experiments are conducted on a Nvidia Tesla A100 GPU. Please refer to Appendix C.3 for more details.
@@ 6 Results
The results are shown in Table 1. It is worth noting that our experiments encompass the use of various LLMs and multiple instances of both exemplar codes and rewritten codes for conducting ablation studies. Here we report the best performance when equipped with ReCo. Comprehensive results can be found in Appendix E. We can observe that ReCo significantly outperforms GAR on both supervised and unsupervised models across diverse search scenarios. With ReCo, the non-neural model BM25 can have competitive performance compared to neural models under zero-shot setting. And ReCo could boost the performance of zero-shot neural models similar to supervised models. We also evaluate ReCo on Contriever, a passage retrieval model that is not specifically trained for code-related tasks. We argue that ReCo can also benefit general-purpose retrieval models. Note that compared with GAR, ReCo does not bring any additional computation in real-time search because the rewritten code could be pre-processed and stored in the codebase.
@@
For the first condition, we analyze the results from BM25 on the four datasets with different LLMs including GPT3.5, Code Llama-7B, 13B, and 34B. Here we do not take the results from dense retrieval systems because neural models can capture the potential relationship among similar tokens. The numerical scores under different evaluation metrics are shown in Appendix E. Regression plots are shown in Fig. 3. According to our first condition, $\Delta\mathrm{MetricScore}$ and $\Delta\mathrm{MRR}$ should be consistent, which means that points are expected to be scattered on Quadrant I and III. We can see that most of the points in CodeBLEU, ROUGE-L, and BLEU are scattered on Quadrant IV. In other words, when the rewritten codes are considered to be more similar to the exemplar code by these metrics, the performance of ReCo, on the contrary, drops compared with GAR. Points from Code Style Similarity mostly fall on Quadrant I and III and the regression line nearly passes through the origin.
For the second condition, we analyze the results from BM25 equipped with GAR. For each query, we calculate the metric score between its exemplar codes and the true code in the codebase and then select the one with the highest metric score. The performance on four datasets after selecting the best exemplar code generated by Code Llama-7B is shown in Table 2. We can observe that compared with random selection, the improvement brought by CodeBLEU, ROUGE-L, and BLEU is not significant generally. On the contrary, Code Style Similarity outperforms other settings. To better understand the preference of different evaluation metrics, we also conduct case studies in Appendix D.
@@
We explore the effect of using different LLMs in ReCo. The results on BM25 are shown in Table 4. The full results including other retrieval models are in Appendix E. GPT3.5’s number of parameters is not released publicly but is estimated at around 175 billion. Generally, we observe that larger LLMs yield greater improvements. However, a decrement in performance is noted with the application of Code Llama-34B. This decrement is attributed to the model’s propensity to generate code not only for the prompted fifth example but also for the initial four in-context examples. Consequently, the generated code is often truncated due to output length limitations.
@@
We also explore the effect of the numbers of generated exemplar codes and rewritten codes in ReCo. The outcomes of BM25 on CoNaLa and MBPP are depicted in Fig. 4, while a comprehensive compilation of results, inclusive of other retrieval models and datasets, can be found in Appendix E. Our observations indicate a marginal enhancement when LLMs are tasked with generating more codes. We discern that the multiple codes generated exhibit similarities, with minor variations, attributable to the self-consistent style of each LLM. To address this, our future work will investigate controlled prompts that steer LLMs towards generating code with controlled stylistic variations, thereby enhancing the diversity of code generation. Fig. 4 also illustrates that the improvement tends to diminish as the quantity of generated codes increases. In practical applications, it is essential to weigh the trade-off between performance enhancement and the incremental costs associated with generation.
7 Discussion
Broader Impact
@@ Conclusion
Appendix A Extended Related Works
Large Language Models
Appendix B Complete Prompt
@@
Appendix C Experiment Settings
C.1 Dataset Statistics
@@ C.2 MRR Calculation
C.3 Implementation Details
For neural models, they all use the same set of hyperparameters. The maximum input length of codes and queries are both set to be 256. Models are trained by Adam and learning rate is set to 5e-6. We adopt mean pooling to get the representation of the whole input sentence to make sure the pooling mechanism is consistent with that during pre-training. The representations are normalized by the L2 norm. UniXcoder is initialized using the publicly available checkpoint at https://huggingface.co/microsoft/unixcoder-base and Contriever is initialized using https://huggingface.co/facebook/contriever-msmarco. CodeBERT is initialized using https://huggingface.co/microsoft/codebert-base and then pre-trained on the CodeSearchNet dataset Husain et al. (2019) for 10 epochs. The pre-training setting is the same as in fine-tuning. All the experiments involving model training are running with 3 random seeds 1234, 12345, and 123456 and they all meet $p<0.01$ of significance tests.
@@ C.3 Implementation Details
Appendix D Case Study
@@
Appendix E Full Results
In this section, we report the full experimental results. The results of $\mathrm{MetricScore}(c_{q},c_{c})$ and $\mathrm{MetricScore}(c_{q},c)$ under different evaluation metrics are shown in Table 6. The results of BM25 are shown in Table 7. The results of fine-tuned UniXcoder and CodeBERT are shown in Table 8 and Table 9, respectively. The results of UniXcoder and Contriever under zero-shot setting are shown in Table 10.
@@ Appendix E Full Results
Table 10: Full results of UniXcoder and Contriever under zero-shot setting. #gen denotes the number of generated and rewritten codes. Bold and underlined results are the best performance of ReCo and the performance of GAR under the same setting, which are reported in Table 1.
@@
Table 11: A prompt example used for generating exemplar codes for the MBPP dataset.
@@
Table 12: A prompt example used for summarizing the original codes for the MBPP dataset.

### v2

@@ Abstract
In code search, the Generation-Augmented Retrieval (GAR) framework, which generates exemplar code snippets to augment queries, has emerged as a promising strategy to address the principal challenge of modality misalignment between code snippets and natural language queries, particularly with the demonstrated code generation capabilities of Large Language Models (LLMs). Nevertheless, our preliminary investigations indicate that the improvements conferred by such an LLM-augmented framework are somewhat constrained. This limitation could potentially be ascribed to the fact that the generated codes, albeit functionally accurate, frequently display a pronounced stylistic deviation from the ground truth code in the codebase. In this paper, we extend the foundational GAR framework and propose a simple yet effective method that additionally Rewrites the Code (ReCo) within the codebase for style normalization. Experimental results demonstrate that ReCo significantly boosts retrieval accuracy across sparse (up to 35.7%), zero-shot dense (up to 27.6%), and fine-tuned dense (up to 23.6%) retrieval settings in diverse search scenarios. To further elucidate the advantages of ReCo and stimulate research in code style normalization, we introduce Code Style Similarity, the first metric tailored to quantify stylistic similarities in code. Notably, our empirical findings reveal the inadequacy of existing metrics in capturing stylistic nuances. The source code and data are available at https://github.com/Alex-HaochenLi/ReCo.
@@ 2 Related Works
Recently, transformer-based pre-trained models significantly outperformed previous methods, since they can be trained on large-scale unlabelled corpus with self-supervised pre-training tasks. Many novel pre-training tasks are proposed to let models have a better general understanding of codes Guo et al. (2021); Li et al. (2022b, c); Shi et al. (2022). For instance, CodeBERT Feng et al. (2020) utilizes masked language modeling and replaced token detection. CodeT5 Wang et al. (2021) focuses on generative tasks through bimodal dual generation. UniXcoder Guo et al. (2022) integrates the aforementioned generative and understanding pre-training tasks. CodeT5+ Wang et al. (2023b) employs the same architecture as CodeT5 and pre-trains it with span denoising, causal language modeling, contrastive learning, and text-code matching from both unimodal code data and bimodal code-text data.
Large Language Models
@@
While LLMs are designed for token generation, their direct application to retrieval tasks such as code search is not suitable. Indeed, there have been attempts to amalgamate the search query and all the candidates together as input, subsequently requesting the LLMs to rank the candidates within the input Qin et al. (2023). However, the constraint on input sequence length impedes its applicability to large-scale retrieval tasks.
One indirect way is to ask LLMs to generate some references and expand the search query with them. This framework, denoted as Generation-Augmented Retrieval, has been proven effective in both question answering and passage retrieval Mao et al. (2020); Gao et al. (2022); Wang et al. (2023a). Mao et al. (2020) is the first work to propose GAR in question answering. HyDE Gao et al. (2022) evaluates GAR in passage retrieval under zero-shot setting. query2doc Wang et al. (2023a) extends GAR to fine-tuning. Our research findings suggest that the Generation-Augmented Retrieval (GAR) method does not substantially enhance the efficiency of code search, primarily due to the significant stylistic difference between exemplar code and true code.
@@ 3.2 ReCo
Detailed examples are provided in Appendix A to further elucidate the prompt.
@@
We offer theoretical insights to differentiate GAR and our proposed ReCo. Each code in the codebase is an implementation of a specific query, which we denote as $c\sim P(q)$. Here $P$ denotes a real-world distribution between queries and codes. LLM also defines a probability distribution over queries. Thus, exemplar codes can be considered to follow $c_{q}\sim\mathrm{LLM}(q)$. We could find that $c$ and $c_{q}$ are sampled from two different distributions given $q$. This accounts for the occasional divergence between the true code and the exemplar code for the same query, as illustrated in Fig. 1.
@@ 5 Experimental Setups
We evaluate ReCo across various search scenarios and programming languages: online forum StackOverflow CoNaLa Yin et al. (2018), coding challenge competence APPS Hendrycks et al. (2021), general programming problems MBPP Austin et al. (2021) and MBJP Athiwaratkun et al. (2022). The first three datasets are written in Python while the last one is written in Java. The statistics of the datasets are shown in Appendix B.1. We take the widely adopted Mean Reciprocal Rank (MRR) as the evaluation metric Li et al. (2023). MRR is the average of reciprocal ranks of the true code snippets for the given query.
@@
We apply ReCo on several models: BM25, an enhanced version of TF-IDF, is a statistical measure that matches certain keywords in codes with the given query. CodeBERT is a bi-modal model pre-trained on Masked Language Modeling and Replaced Token Detection. UniXcoder unifies both generating and understanding pre-training tasks to further enhance code representation learning by leveraging cross-modal contents. Contriever Izacard et al. (2021) is an unsupervised dense information retrieval model that leverages contrastive learning for its training. CodeT5+ is pre-trained on both unimodal code data and bimodal code-text data with a diverse set of pretraining tasks including span denoising, causal language modeling, contrastive learning, and text-code matching.
@@
CodeT5+ 73.7 59.4 7.6 67.7
   +  GAR 80.3 77.7 10.2 79.2
   +  ReCo $\textbf{80.8}_{+0.5}$ $\textbf{89.4}_{+11.7}$ $\textbf{29.9}_{+19.7}$ $\textbf{84.0}_{+4.8}$
@@
Compared Metrics
Implementation Details
When prompting LLMs, we randomly sample 4 in-context examples from the training sets and set a temperature of 1. And when we prompt LLMs multiple times for the same input, each time we resample the in-context example. The maximum length of output for code summarization and generation is 128 and 256, respectively. For sparse retrieval, we use the default implementation from Pyserini Lin et al. (2021). For dense retrieval, during training, we adopt the default hyperparameters described in the original paper. They are trained for 10 epochs with a batch size of 32. Experiments are conducted on a Nvidia Tesla A100 GPU. Please refer to Appendix B.3 for more details.
@@ 6 Results
The results are shown in Table 1. It is worth noting that our experiments encompass the use of various LLMs and multiple instances of both exemplar codes and rewritten codes for conducting ablation studies. Here we report the best performance when equipped with ReCo. Comprehensive results can be found in Appendix D. We can observe that ReCo significantly outperforms GAR on both supervised and unsupervised models across diverse search scenarios. With ReCo, the non-neural model BM25 can have competitive performance compared to neural models under zero-shot setting. And ReCo could boost the performance of zero-shot neural models similar to supervised models. We also evaluate ReCo on Contriever, a passage retrieval model that is not specifically trained for code-related tasks. We argue that ReCo can also benefit general-purpose retrieval models. Note that compared with GAR, ReCo does not bring any additional computation in real-time search because the rewritten code could be pre-processed and stored in the codebase.
@@
For the first condition, we analyze the results from BM25 on the four datasets with different LLMs including GPT3.5, Code Llama-7B, 13B, and 34B. Here we do not take the results from dense retrieval systems because neural models can capture the potential relationship among similar tokens. The numerical scores under different evaluation metrics are shown in Appendix D. Regression plots are shown in Fig. 3. According to our first condition, $\Delta\mathrm{MetricScore}$ and $\Delta\mathrm{MRR}$ should be consistent, which means that points are expected to be scattered on Quadrant I and III. We can see that most of the points in CodeBLEU, ROUGE-L, and BLEU are scattered on Quadrant IV. In other words, when the rewritten codes are considered to be more similar to the exemplar code by these metrics, the performance of ReCo, on the contrary, drops compared with GAR. Points from Code Style Similarity mostly fall on Quadrant I and III and the regression line nearly passes through the origin.
For the second condition, we analyze the results from BM25 equipped with GAR. For each query, we calculate the metric score between its exemplar codes and the true code in the codebase and then select the one with the highest metric score. The performance on four datasets after selecting the best exemplar code generated by Code Llama-7B is shown in Table 2. We can observe that compared with random selection, the improvement brought by CodeBLEU, ROUGE-L, and BLEU is not significant generally. On the contrary, Code Style Similarity outperforms other settings. To better understand the preference of different evaluation metrics, we also conduct case studies in Appendix C.
@@
We explore the effect of using different LLMs in ReCo. The results on BM25 are shown in Table 4. The full results including other retrieval models are in Appendix D. GPT3.5’s number of parameters is not released publicly but is estimated at around 175 billion. Generally, we observe that larger LLMs yield greater improvements. However, a decrement in performance is noted with the application of Code Llama-34B. This decrement is attributed to the model’s propensity to generate code not only for the prompted fifth example but also for the initial four in-context examples. Consequently, the generated code is often truncated due to output length limitations.
@@
We also explore the effect of the numbers of generated exemplar codes and rewritten codes in ReCo. The outcomes of BM25 on CoNaLa and MBPP are depicted in Fig. 4, while a comprehensive compilation of results, inclusive of other retrieval models and datasets, can be found in Appendix D. Our observations indicate a marginal enhancement when LLMs are tasked with generating more codes. We discern that the multiple codes generated exhibit similarities, with minor variations, attributable to the self-consistent style of each LLM. To address this, our future work will investigate controlled prompts that steer LLMs towards generating code with controlled stylistic variations, thereby enhancing the diversity of code generation. Fig. 4 also illustrates that the improvement tends to diminish as the quantity of generated codes increases. In practical applications, it is essential to weigh the trade-off between performance enhancement and the incremental costs associated with generation.
7 Broader Impact
@@ 8 Conclusion
Acknowledgement
We thank the anonymous reviewers for their helpful comments and suggestions.
Limitations
There are mainly two limitations of this work. First, although ReCo does not require any additional computation in real-time search compared with GAR, both GAR and ReCo rely on the real-time generation of exemplar codes. Therefore, ReCo and GAR may have limitations when applied to tasks that demand low latency. The latency of generating exemplar codes depends on the time cost of LLM inference. As stated in research works focusing on GAR, over the years the cost of hardware has decreased a lot and there are many works proposed to improve the inference efficiency of LLMs Gao et al. (2022); Wang et al. (2023a). We believe the efficiency problem of GAR and ReCo will be addressed in the future. The second limitation is that we do not evaluate ReCo on some extremely large-scale codebases like CodeSearchNet Husain et al. (2019). This is due to the time burden of generating exemplar codes and rewriting codes. For example, according to our estimation, there are 1,005,474 queries in total in CodeSearchNet hence generating one exemplar code for them costs more than two months. To address this limitation, we evaluate ReCo on several search scenarios covering coding challenge competence, online programming community, and general programming problems to show the effectiveness of ReCo.
Appendix A Complete Prompt
@@
Appendix B Experiment Settings
B.1 Dataset Statistics
@@ B.2 MRR Calculation
B.3 Implementation Details
For neural models, they all use the same set of hyperparameters. The maximum input length of codes and queries are both set to be 256. Models are trained by Adam and learning rate is set to 5e-6. We adopt mean pooling to get the representation of the whole input sentence to make sure the pooling mechanism is consistent with that during pre-training. The representations are normalized by the L2 norm. UniXcoder is initialized using the publicly available checkpoint at https://huggingface.co/microsoft/unixcoder-base, Contriever is initialized using https://huggingface.co/facebook/contriever-msmarco, and CodeT5+ is initialized using https://huggingface.co/Salesforce/codet5p-110m-embedding. CodeBERT is initialized using https://huggingface.co/microsoft/codebert-base and then pre-trained on the CodeSearchNet dataset Husain et al. (2019) for 10 epochs. The pre-training setting is the same as in fine-tuning. All the experiments involving model training are running with 3 random seeds 1234, 12345, and 123456 and they all meet $p<0.01$ of significance tests.
Appendix C Case Study
@@
Appendix D Full Results
In this section, we report the full experimental results. The results of $\mathrm{MetricScore}(c_{q},c_{c})$ and $\mathrm{MetricScore}(c_{q},c)$ under different evaluation metrics are shown in Table 6. The results of BM25 are shown in Table 7. The results of fine-tuned UniXcoder and CodeBERT are shown in Table 8 and Table 9, respectively. The results of UniXcoder, Contriever, and CodeT5+ under zero-shot setting are shown in Table 10.
@@ Appendix D Full Results
[…]
ReCo 57.7 67.2 29.9 73.0
[…]
ReCo 72.8 89.4 24.7 84.0
CodeLlama-34B 4 GAR 74.5 68.4 8.1 76.5
ReCo 67.6 57.9 15.8 62.9
[…]
ReCo 80.8 82.2 19.1 81.4
Table 10: Full results of UniXcoder, Contriever, and CodeT5+ under zero-shot setting. #gen denotes the number of generated and rewritten codes. Bold and underlined results are the best performance of ReCo and the performance of GAR under the same setting, which are reported in Table 1.
@@
Table 11: A prompt example used for generating exemplar codes for the MBPP dataset. A more detailed prompt may increase the quality of the exemplar code and we leave this as our future work.
@@
Table 12: A prompt example used for summarizing the original codes for the MBPP dataset. A more detailed prompt may increase the quality of the rewritten code and we leave this as our future work.